OpenStudy (anonymous):

1. Solve for x in the interval [0 , 2 π ). (a) 2sin^2x=sin x (b) 3sin^2x+cos^2x=2 (c) 1+3cosx=cos2x (d) sin2x+sin x=0

OpenStudy (anonymous):

a) \(2\sin^2x-\sin x=0 \implies \sin x(2\sin{x}-1)=0 \implies \sin x=0 \text{ or } \sin{x}=\frac{1}{2}.\) You can proceed form here.

OpenStudy (anonymous):

dont know hot to solve it